John is 7 times as old as his son Ron now. In two years, he will only be 5 times the age of his sone then. How old are they now?
Answer by JulietG(1812)

J = 7R [John is 7 times as old as his son Ron now.]
J + 2 = 5(R+2) [In two years, he will only be 5 times the age of his son then]
Substitute the value of J from the first equation into the second.
(7r) + 2 = 5(R+2)
Distribute
7R + 2 = 5R + 10
Subtract 5R from each side
2R + 2 = 10
Subtract 2 from each side
2R = 8
Divide each side by 2
R = 4, which means J = 28
.
Let's plug the values in to prove.
In two years, R will be 6 and J will b 30
Is 30 five time 6? Yes.
Success!